### v2

@@ 4.6 Local observables
$\langle\Phi_{1}|\Phi^{\textrm{(cyl.)}}_{2}(\hat{\mathbf{n}},\tau)\Phi^{\textrm{(cyl.)}}_{3}(\hat{\mathbf{z}})|\Phi_{4}\rangle=\frac{e^{\Delta_{\Phi_{2}}\tau/R}}{R^{\Delta_{\Phi_{2}}+\Delta_{\Phi_{3}}}}\langle\Phi_{1}^{\dagger}(\infty)\Phi_{2}(\mathbf{r})\Phi_{3}(\hat{\mathbf{z}})\Phi_{4}(0)\rangle,$ (54)
@@
$\langle 0|\Phi^{\textrm{(cyl.)}}_{2}(\hat{\mathbf{n}})\Phi^{\textrm{(cyl.)}}_{2}(\hat{\mathbf{z}})|0\rangle=R^{-2\Delta_{\Phi_{2}}}\langle\Phi_{2}(\hat{\mathbf{n}})\Phi_{2}(\hat{\mathbf{z}})\rangle\\ =\frac{1}{R^{2\Delta_{\Phi_{2}}}|\hat{\mathbf{n}}-\hat{\mathbf{z}}|^{2\Delta_{\Phi_{2}}}}=\frac{1}{R^{2\Delta_{\Phi_{2}}}(1-\cos\theta)^{\Delta_{\Phi_{2}}}}.$ (56)
@@ 4.7 Conformal generators
A general Noether current and corresponding generator of the infinitesimal spacetime transformation $x^{\mu}\mapsto x^{\mu}+\epsilon^{\mu}(x)$ can be expressed in terms of the stress tensor
@@
$P^{\mu}$ $=\int\mathrm{d}^{2}\hat{\mathbf{n}}\,(n^{\mu}T^{0}{}_{0}+iT^{0\mu}),$
$K^{\mu}$ $=\int\mathrm{d}^{2}\hat{\mathbf{n}}\,(n^{\mu}T^{0}{}_{0}-iT^{0\mu}).$ (58)
@@
$\Lambda_{m}=P_{m}+K_{m}=\sqrt{\frac{16\pi}{3}}\int\mathrm{d}^{2}\hat{\mathbf{n}}\,\bar{Y}_{1m}(\hat{\mathbf{n}})\mathscr{H}(\hat{\mathbf{n}}).$ (59)
@@
[…]
The derivation of the expression and the constant factors are calculated and given in Ref. [14, 15].
We then need to find the expression for the Hamiltonian density. For example, for the Ising model, it is the local density operator and density-density interactions with some full derivatives
$\mathscr{H}(\hat{\mathbf{n}})=n_{z}\left(g_{0}+g_{1}\nabla^{2}\right)n_{z}-hn_{x}+g_{D,1}\nabla^{2}n_{x}+g_{D,2}\nabla^{2}n_{z}^{2}+\dots,$ (61)
where $g_{D,i}$ are undetermined constants that does not affect the Hamiltonian $H=\int\mathrm{d}^{2}\hat{\mathbf{n}}\,\mathscr{H}$. We have only listed a few examples of the allowed full derivatives.
@@
$\sum_{\begin{subarray}{c}l_{1}l_{2}m_{1}m_{2}\end{subarray}}\tilde{U}_{l_{1}l_{2}}\Delta^{\dagger}_{l_{1}m_{1}}\Delta_{l_{2}m_{2}}\langle(2s-l_{1})(-m_{1}),(2s-l_{2})m_{2}|1m\rangle$ (62)
For $l_{1}\in 2\mathbb{Z}$, the spin index in the pairing operator is antisymmetrised ; For $l_{1}\in 2\mathbb{Z}+1$, the spin index in the pairing operator is symmetrised. Therefore, $l_{1}-l_{2}\in 2\mathbb{Z}$ for non-zero results. And since $|l_{1}-l_{2}|\leq 1$, we conclude $l_{1}=l_{2}$. so
@@
After obtaining $\Lambda^{\mu}=P^{\mu}+K^{\mu}$, the separate $P^{\mu}$ and $K^{\mu}$ can be obtained by considering the commutator with the dilatation generator $D$, which is proportional to the Hamiltonian. As $[D,P^{\mu}]=P^{\mu}$ and $[D,K^{\mu}]=-K^{\mu}$.
@@ 5.1 Exact diagonalisation (ED)
The ED calculation can be optimised in several ways. The storage of the Hamiltonian matrix may be compressed by data structure tailored for sparse matrix such as compressed sparse column (CSC). The Hamiltonian matrix is usually block diagonal due to the symmetry of the Hamiltonian. The Hilbert space is divided into several sectors that carry different representations under the symmetry, and acting the Hamiltonian on a state in a sector results in a state in the same sector. E.g., in the ED calculation for the Ising model on the fuzzy sphere, the symmetries we can use include two $\mathrm{U}(1)$ symmetries, viz. the conservation of particle number and the angular momentum in the $z$-direction, and three $\mathbb{Z}_{2}$ symmetries, viz. the Ising $\mathbb{Z}_{2}$ flavour symmetry, the particle-hole symmetry and the $\pi$-rotation along the $y$-axis [^11] So far, FuzzifiED only supports $\mathrm{U}(1)$ and $\mathbb{Z}_{p}$ symmetries..
The ED method enjoys several advantages, including (1) the full knowledge of the eigenstate wavefunction and (2) the ability to access relatively high excited states. However, despite these optimisations, the dimension of the Hilbert space scales exponentially with the number of orbitals. This results in exponentially growing space and time complexity. E.g., for the Ising model on the fuzzy sphere, for $N_{m}=14$, the dimension of Hilbert space $\dim\mathscr{H}=1.8\times 10^{5}$ and the number of elements in the Hamiltonian is $N_{\textrm{el}}=1.1\times 10^{7}$ ; for $N_{m}=16$, the numbers have already grown to $\dim\mathscr{H}=2.2\times 10^{6}$ and $N_{\textrm{el}}=2.1\times 10^{8}$, which translates to a memory demand of $3.1$ gigabytes.
@@ 5.2 Density matrix renormalisation group (DMRG)
To overcome the size limit of ED, the density matrix renormalisation group (DMRG) is a powerful method to calculate the ground state of a quasi-one-dimensional system. It was first invented by White [104] as an improvement to the numerical renormalisation group (NRG) used in the Kondo problem. Since its proposal, it has been proven potent in various problems in condensed matter physics, such as the static and dynamic properties of one-dimensional models such as the Heisenberg, $t$$J$ and Hubbard models [105]. Later, Schollwöck has discovered a new point of view that implements the DMRG in the language of matrix product states (MPS) [106].
Briefly speaking, in this language, DMRG is a variational method that optimises the fidelity between the exact ground state and the variational MPS. During each ‘sweep’, DMRG solves a local eigenvalue problem for the active tensors to improve the approximation. To find the excited states, one needs to add projection $|0\rangle\langle 0|$ of the ground state $|0\rangle$ to the Hamiltonian by hand.
Although the fuzzy sphere deals with $(2+1)$-dimensional quantum systems, the basis of the lowest Landau level provides a natural way to express it as a quasi-1d problem. Therefore, DMRG has been a powerful numerical method for the fuzzy sphere. However, like other $(2+1)$d models, the DMRG on the fuzzy sphere also suffers from the divergence of the required maximal bond dimension with system size. One should thus be careful with the convergence of the results when doing DMRG.
@@ 6 Installation and usage
The package FuzzifiED is implemented in the Julia language [21] [^12] The most intensive core functionalities are written in Fortran and wrapped by Julia interfaces, but the users are not required to have any interaction with or knowledge of the Fortran part of the source code.. Some useful links are given in Table 1. To install the package, enter julia to enter the Julia REPL (read-eval-print loop), and then run the following command.
@@
To obtain the documentation for an interface, type ‘?’ followed by the keyword in the Julia REPL, e.g., ‘?Confs’.
@@ 7 Exact diagonalisation
7.1 Setup
Before starting the calculation, we need to input the setup for the system, including the number of flavours $N_{f}$, orbitals $N_{m}$ and sites $N_{o}$. Here we specify some of our notations.
  • A ‘flavour’ is labelled by $f$. The number of flavours is $N_{f}$.
  • An ‘orbital’ is specified by the magnetic quantum number labelled by $m$. The number of orbitals is $N_{m}=2s+1$.
  • A ‘site’ is specified by both the flavour and the orbital index $o=(f,m)$. The number of sites is $N_{o}=N_{m}N_{f}$. In practice, we label the sites with an integer from $1$ to $N_{o}$. We store the sites in an ascending order of first $m$ and then $f$ : $o=(m+s)N_{f}+f$.
@@ 7.2 Constructing the configurations
The first step for the ED calculation is to construct the basis that respects the symmetries of the Hamiltonian. This is divided into two procedures : (1) to generate the ‘configurations’ that carry the diagonal quantum numbers, and (2) to generate the ‘basis’ that also carries the off-diagonal quantum numbers (under discrete transformations). The ‘configurations’ are the collection of states that can be written as direct products of occupied $|1\rangle$ or empty $|0\rangle$ on each site and carry certain diagonal quantum numbers (QNDiag).
@@
  QNDiag — Type  The type contains the fields
  • name :: String — the name of the diagonal quantum number [^13] The name is only needed for conversion into quantum numbers in ITensor..
@@
  • GetFlavQNDiag(nm, nf, qf[, id, modul]) — a linear combination of number of electrons in each flavour $Q=\sum_{f}q_{f}n_{f}$, where $\{q_{f}\}$ is stored in qf in the format of either an array or a dictionary [^15] E.g., for $Q=n_{f=1}-n_{f=3}$ in a 4-flavour system, both qf = [1, 0, -1, 0] and qf = Dict(1 => 1, 3 => -1) are acceptable..
@@
  Confs — Type  The mutable type contains the fields
@@
where qnd is the array of QNDiags, and secd is the array of charges $Q$ of each QNDiag [^16] In general, many methods in FuzzifiED admits keyword arguments num_th :: Int64 that specifies the number of threads and disp_std :: Bool that specifies whether or not the log shall be displayed. Hereafter, we will omit these two arguments..
Here, each configuration is stored as a binary number with $N_{o}$ bits. If the $o$-th site in the configuration is occupied, the $(o-1)$-th bit of the configuration is $1$; if the site is empty, then the bit is 00. Besides the storage of the configuration, we also need a reverse look-up process that returns the index from the binary string. This is realised by a Lin table stored in lid and rid. The details are given in Appendix A.1.
@@ 7.3 Constructing the basis
Having constructed the configurations, we now construct the basis of the Hilbert space. The ‘basis’ is the collection of states that are linear combinations of the configurations carrying certain diagonal and $\mathbb{Z}_{p}$ off-diagonal quantum numbers (QNOffd).
@@
  QNOffd — Type  The mutable type contains the fields
@@
  Basis — Type  The mutable type contains the fields
@@ 7.4 Recording the many-body operator terms
  Term — Type  The mutable type contains the fields
@@
[…]
@@ 7.5 Generating sparse matrix
  Operator — Type  The type can be initialised with the method
@@
  OpMat{T} — Type  where T is the type of the elements, it can either be ComplexF64 or Float64. It contains the fields
@@ 7.6 Finding eigenstates
  • A dimd×$\times$×nst matrix where every column records an eigenstate.
@@ 7.7 Inner product of states, operators and transformations
where T is the type of the elements. For example, the code to measure the angular momenta of each state is
@@
  Transf — Type  The type can be initialised from a QNOffd by
@@ 7.8 Measuring local observables
$\mathscr{O}(\hat{\mathbf{n}})$ $=\sum_{lm}Y^{(s)}_{lm}(\hat{\mathbf{n}})\mathscr{O}_{lm}$
@@
×$\qquad\qquad\times(-1)^{s+m}\sqrt{\frac{(2l_{1}+1)(2l_{2}+1)(2l_{3}+1)}{4\pi}}\begin{pmatrix}l_{1}&l_{2}&l\\ m_{1}&m_{2}&-m\end{pmatrix}\begin{pmatrix}l_{1}&l_{2}&l\\ -s_{1}&-s_{2}&s\end{pmatrix}$ (78)
@@
  SphereObs — Type  The type contains the fields
@@
  • StoreComps(obs) calculates and stores each component of the observable.
  • Laplacian(obs) takes the Laplacian of an observable.
  • GetComponent(obs, l, m) returns a spherical component of the observable $\mathscr{O}_{lm}$.
  • GetPointValue(obs, theta, phi) returns an observable at one point $\mathscr{O}(\hat{\mathbf{n}})$.
  • GetIntegral(obs) returns the uniform spatial integral of an observable.
  • FilterComponent(obs, flt) filters a certain set of modes ofan observable.
Several important types of spherical observables are built-in in FuzzifiED [^17] By default, the operators are normalised by angular coordinate. A parameter ObsNormRadSq is used to normalise the observables on a sphere with radius $R$.
  • GetElectronObs(nm, nf, f) — electron annihilation operator $\psi_{f}(\hat{\mathbf{n}})$.
  • GetDensityObs(nm, nf[, mat]) — density operator $n_{M}(\hat{\mathbf{n}})=\sum_{ff^{\prime}}\psi^{\dagger}_{f}(\hat{\mathbf{n}})M_{ff^{\prime}}\psi_{f^{\prime}}(\hat{\mathbf{n}})$.
  • GetPairingObs(nm, nf, mat) — pair operator $\Delta_{M}(\hat{\mathbf{n}})=\sum_{ff^{\prime}}\psi_{f}(\hat{\mathbf{n}})M_{ff^{\prime}}\psi_{f^{\prime}}$.

### v3

@@ 4.6 Local Observables
$\langle\Phi_{1}|\Phi^{\textrm{(cyl.)}}_{2}(\mathbf{r},\tau)\Phi^{\textrm{(cyl.)}}_{3}(\hat{\mathbf{z}})|\Phi_{4}\rangle=\frac{e^{\Delta_{\Phi_{2}}\tau/R}}{R^{\Delta_{\Phi_{2}}+\Delta_{\Phi_{3}}}}\langle\Phi_{1}^{\dagger}(\infty)\Phi_{2}(x)\Phi_{3}(\hat{\mathbf{z}})\Phi_{4}(0)\rangle,$ (53)
@@
$\langle 0|\Phi^{\textrm{(cyl.)}}_{2}(\mathbf{r})\Phi^{\textrm{(cyl.)}}_{2}(\hat{\mathbf{z}})|0\rangle=R^{-2\Delta_{\Phi_{2}}}\langle\Phi_{2}(\mathbf{r})\Phi_{2}(\hat{\mathbf{z}})\rangle\\ =\frac{1}{R^{2\Delta_{\Phi_{2}}}|\mathbf{r}-\hat{\mathbf{z}}|^{2\Delta_{\Phi_{2}}}}=\frac{1}{R^{2\Delta_{\Phi_{2}}}(1-\cos\theta)^{\Delta_{\Phi_{2}}}}.$ (55)
@@ 4.7 Conformal Generators
A general Noether current and corresponding generator of the infinitesimal space-time transformation $x^{\mu}\mapsto x^{\mu}+\epsilon^{\mu}(x)$ can be expressed in terms of the stress tensor
@@
$P^{\mu}$ $=\int\mathrm{d}^{2}\mathbf{r}\,(r^{\mu}T^{0}{}_{0}+iT^{0\mu}),$
$K^{\mu}$ $=\int\mathrm{d}^{2}\mathbf{r}\,(r^{\mu}T^{0}{}_{0}-iT^{0\mu}).$ (57)
@@
$\Lambda_{m}=P_{m}+K_{m}=\sqrt{\frac{16\pi}{3}}\int\mathrm{d}^{2}\mathbf{r}\,\bar{Y}_{1m}(\hat{\mathbf{n}})\mathscr{H}(\mathbf{r}).$ (58)
@@
$\Lambda^{\mu}|\Phi\rangle$ $=\textrm{const.}\times|\partial^{\mu}\Phi\rangle$
$\Lambda^{\mu}|\partial_{\mu}\Phi\rangle$ $=\textrm{const.}\times|\Phi\rangle+\textrm{const.}\times|\partial^{\mu}\partial^{\nu}\Phi\rangle+\textrm{const.}\times|\Box\Phi\rangle.$ (59)
The derivation of the expression and the constant factors are calculated and given in Refs. [14, 15].
We then need to find the expression for the Hamiltonian density. E. g., for the Ising model, it is the local density operator and density-density interactions with some full derivatives
$\mathscr{H}(\mathbf{r})=n_{z}\left(g_{0}+g_{1}\nabla^{2}\right)n_{z}-hn_{x}+g_{D,1}\nabla^{2}n_{x}+g_{D,2}\nabla^{2}n_{z}^{2}+\dots,$ (60)
where $g_{D,i}$ are undetermined constants that does not affect the Hamiltonian $H=\int\mathrm{d}^{2}\mathbf{r}\,\mathscr{H}$. We have only listed a few examples of the allowed full derivatives.
@@
$\sum_{\begin{subarray}{c}l_{1}l_{2}m_{1}m_{2}\end{subarray}}\tilde{U}_{l_{1}l_{2}}\Delta^{\dagger}_{l_{1}m_{1}}\Delta_{l_{2}m_{2}}\langle(2s-l_{1})m_{1},(2s-l_{2})(-m_{2})|1m\rangle$ (61)
For $l_{1}\in 2\mathbb{Z}$, the spin index in the pairing operator is anti-symmetrised ; For $l_{1}\in 2\mathbb{Z}+1$, the spin index in the pairing operator is symmetrised. Therefore, $l_{1}-l_{2}\in 2\mathbb{Z}$ for non-zero results. And since $|l_{1}-l_{2}|\leq 1$, we conclude $l_{1}=l_{2}$. so
@@
After obtaining $\Lambda^{\mu}=P^{\mu}+K^{\mu}$, the separate $P^{\mu}$ and $K^{\mu}$ can be obtained by considering the commutator with the dilatation generator $D$, which is proportional to the Hamiltonian. As $D,P^{\mu}]=P^{\mu}$ and $D,K^{\mu}]=-K^{\mu}$,
@@ 5.1 Exact Diagonalisation (ED)
The ED calculation can be optimised in several ways. The storage of the Hamiltonian matrix may be compressed by data structure tailored for sparse matrix such as compressed sparse column (CSC). The Hamiltonian matrix is usually block diagonal due to the symmetry of the Hamiltonian. The Hilbert space is divided into several sectors that carry different representations under the symmetry, and acting the Hamiltonian on a state in a sector results in a state in the same sector. E. g., in the ED calculation for the Ising model on the fuzzy sphere, the symmetries we can use include two $\mathrm{U}(1)$ symmetries, viz. the conservation of particle number and the angular momentum in the $z$-direction, and three $\mathbb{Z}_{2}$ symmetries, viz. the Ising $\mathbb{Z}_{2}$ flavour symmetry, the particle-hole symmetry and the $\pi$-rotation along the $y$-axis [^12] So far, FuzzifiED only supports $\mathrm{U}(1)$ and $\mathbb{Z}_{p}$ symmetries.
The ED method enjoys several advantages, including (1) the full knowledge of the eigenstate wave-function and (2) the ability to access relatively high excited states. However, despite these optimisations, the dimension of the Hilbert space scales exponentially with the number of orbitals. This results in exponentially growing space and time complexity. E. g., for the Ising model on the fuzzy sphere, for $N_{m}=14$, the dimension of Hilbert space $\dim\mathscr{H}=1.8\times 10^{5}$ and the number of elements in the Hamiltonian is $N_{\text{el}}=1.1\times 10^{7}$ ; for $N_{m}=16$, the numbers have already grown to $\dim\mathscr{H}=2.2\times 10^{6}$ and $N_{\text{el}}=2.1\times 10^{8}$, which translates to a memory demand of $3.1$ gigabytes.
@@ 5.2 Density Matrix Renormalisation Group (DMRG)
To overcome the size limit of ED, the density matrix renormalisation group (DMRG) is a powerful method to calculate the ground state of a quasi-one-dimensional system. It was first invented by White [148] as an improvement to the numerical renormalisation group (NRG) used in the Kondo problem. Since its proposal, it has been proven potent in various problems in condensed-matter physics, such as the static and dynamic properties of one-dimensional models such as the Heisenberg, ttJJ and Hubbard models [149]. Later, Schollwöck has discovered a new point of view that implements the DMRG in the language of matrix product states (MPS) [150].
Briefly speaking, in this language, DMRG is a variational method that optimises the fidelity between the exact ground state and the variational MPS. During each “sweep,” DMRG solves a local eigenvalue problem for the active tensors to improve the approximation. To find the excited states, one needs to add projection $|0\rangle\langle 0|$ of the ground state $|0\rangle$ to the Hamiltonian by hand.
Although the fuzzy sphere deals with $(2+1)$-dimensional quantum systems, the basis of the lowest Landau level provides a natural way to express it as a quasi-1D problem. Therefore, DMRG has been a powerful numerical method for the fuzzy sphere. However, like other (2+1)(2+1)D models, the DMRG on the fuzzy sphere also suffers from the divergence of the required maximal bond dimension with system size. One should thus be careful with the convergence of the results when doing DMRG.
@@ 6 Installation and Usage
The package FuzzifiED is implemented in the Julia language [36]. [^13] The most intensive core functionalities are written in Fortran and wrapped by Julia interfaces, but the users are not required to have any interaction with or knowledge of the Fortran part of the source code. Some useful links are given in Table 1. To install the package, enter julia to enter the Julia REPL (read-eval-print loop), and then run the following command.
@@
To obtain the documentation for an interface, type “?” followed by the keyword in the Julia REPL, e. g.?Confs.”
@@ 7 Exact Diagonalisation
7.1 Set-up
Before starting the calculation, we need to input the set-up for the system, including the number of flavours $N_{f}$, orbitals $N_{m}$ and sites $N_{o}$. Here we specify some of our notations.
  • A “flavour” is labelled by $f$. The number of flavours is $N_{f}$.
  • An “orbital” is specified by the magnetic quantum number labelled by $m$. The number of orbitals is $N_{m}=2s+1$.
  • A “site” is specified by both the flavour and the orbital index $o=(f,m)$. The number of sites is $N_{o}=N_{m}N_{f}$. In practice, we label the sites with an integer from $1$ to $N_{o}$. We store the sites in an ascending order of first $m$ and then $f$ : $o=(m+s)N_{f}+f$.
@@ 7.2 Constructing the Configurations
The first step for the ED calculation is to construct the basis that respects the symmetries of the Hamiltonian. This is divided into two procedures : (1) to generate the “configurations” that carry the diagonal quantum numbers, and (2) to generate the “basis” that also carries the off-diagonal quantum numbers (under discrete transformations). The “configurations” are the collection of states that can be written as direct products of occupied $|1\rangle$ or empty $|0\rangle$ on each site and carry certain diagonal quantum numbers (QNDiag).
@@
QNDiag — Type
The type contains the fields
  • name :: String — the name of the diagonal quantum number. [^14] The name is only needed for conversion into quantum numbers in ITensor.
@@
  • GetFlavQNDiag(nm, nf, qf[, id, modul]) — a linear combination of number of electrons in each flavour $Q=\sum_{f}q_{f}n_{f}$, where $\{q_{f}\}$ is stored in qf in the format of either an array or a dictionary. [^16] E. g., for $Q=n_{f=1}-n_{f=3}$ in a 4-flavour system, both qf = [1, 0, -1, 0] and qf = Dict(1 => 1, 3 => -1) are acceptable.
@@
Confs — Type
The mutable type contains the fields
@@
where qnd is the array of QNDiags, and secd is the array of charges $Q$ of each QNDiag. [^17] In general, many methods in FuzzifiED admits keyword arguments num_th :: Int64 that specifies the number of threads and disp_std :: Bool that specifies whether or not the log shall be displayed. Hereafter, we will omit these two arguments.
Here, each configuration is stored as a binary number with $N_{o}$ bits. If the $o$-th site in the configuration is occupied, the (o1)(o-1)-th bit of the configuration is $1$; if the site is empty, then the bit is 0. Besides the storage of the configuration, we also need a reverse look-up process that returns the index from the binary string. This is realised by a Lin table stored in lid and rid. The details are given in Appendix A.1.
@@ 7.3 Constructing the Basis
Having constructed the configurations, we now construct the basis of the Hilbert space. The “basis” is the collection of states that are linear combinations of the configurations carrying certain diagonal and $\mathbb{Z}_{p}$ off-diagonal quantum numbers (QNOffd).
@@
QNOffd — Type
The mutable type contains the fields
@@
Basis — Type
The mutable type contains the fields
@@ 7.4 Recording the Many-Body Operator Terms
Term — Type
The mutable type contains the fields
@@
    $C_{2}=\sum_{imm^{\prime}\{f_{i}\}}\frac{(c^{\dagger}_{mf_{1}}G^{i}_{f_{1}f_{2}}c_{mf_{2}})(c^{\dagger}_{m^{\prime}f_{3}}(G^{i}_{f_{3}f_{4}})^{\dagger}c_{m^{\prime}f_{4}})}{\operatorname{tr}G_{i}^{\dagger}G_{i}}-\textrm{(trace)}$ (74)
@@ 7.5 Generating Sparse Matrix
Operator — Type
The type can be initialised with the method
@@
OpMat{T} — Type
where T is the type of the elements, it can either be ComplexF64 or Float64. It contains the fields
@@ 7.6 Finding Eigenstates
[…]
@@ 7.7 Inner Product of States, Operators and Transformations
where T is the type of the elements. E. g., the code to measure the angular momenta of each state is
@@
Transf — Type
The type can be initialised from a QNOffd by
@@ 7.8 Measuring Local Observables
$\mathscr{O}(\mathbf{r})$ $=\sum_{lm}Y^{(s)}_{lm}(\hat{\mathbf{n}})\mathscr{O}_{lm}$
@@
$\qquad\qquad\times(-1)^{s+m}\sqrt{\frac{(2l_{1}+1)(2l_{2}+1)(2l_{3}+1)}{4\pi}}\begin{pmatrix}l_{1}&l_{2}&l\\ m_{1}&m_{2}&-m\end{pmatrix}\begin{pmatrix}l_{1}&l_{2}&l\\ -s_{1}&-s_{2}&s\end{pmatrix}$ (77)
@@
SphereObs — Type
The type contains the fields
@@
Several important types of spherical observables are built-in in FuzzifiED [^18] Two parameters ObsNormRadSq and ObsMomIncr control how the components are normalised and whether they increase or decrease $L^{z}$.
  • GetElectronObs(nm, nf, f) — electron annihilation operator $\psi_{f}(\mathbf{r})$.
  • GetDensityObs(nm, nf[, mat]) — density operator $n_{M}(\mathbf{r})=\sum_{ff^{\prime}}\psi^{\dagger}_{f}(\mathbf{r})M_{ff^{\prime}}\psi_{f^{\prime}}(\mathbf{r})$.
  • GetPairingObs(nm, nf, mat) — pair operator $\Delta_{M}(\mathbf{r})=\sum_{ff^{\prime}}\psi_{f}(\mathbf{r})M_{ff^{\prime}}\psi_{f^{\prime}}$.